Skip to content

OSG Technology Area Meeting, 18 September 2017

Coordinates: Conference: 719-284-5267, PIN: 57363; https://www.uberconference.com/osgblin
Attending:

Announcements

Triage Duty

  • This week: Suchandra
  • Next week: TimT
  • 10 (+0) open tickets

JIRA

# of tickets Δ State
162 +9 Open
19 +2 In Progress
1 -5 Ready for Testing
0 -20 Ready for Release

Release Schedule

Name Version Development Freeze Package Freeze Release Notes
October 3.4.4, 3.3.29 2017-09-25 2017-10-02 2017-10-10
November 3.4.5, 3.3.30 2017-10-30 2017-11-06 2017-11-14 5 week cycle
December 3.4.6, 3.3.31 2017-11-27 2017-12-04 2017-12-12

Notes: Additional “urgent” releases may be scheduled for the 4th Tuesday of each month. The Testing date is when acceptance testing will be scheduled for releasable packages; if a package is added after this date, it may not be possible to schedule adequate testing time, thereby forcing it into the next release.

OSG Software Team

osghost downtime starts tomorrow at 2pm Central

Documentation

  • https://github.com/opensciencegrid/docs/pulse#new-issues
    • 7 docs fully migrated (many more awaiting review!), >100 docs archived
    • Release3: 17 high-priority and ~30 low-priority docs remaining
    • SoftwareTeam: ~56 docs remaining
  • Carl's mass migration and archival wrapper tools are robust enough to be used by other areas

Discussions

None this week

Support Update

  • Clemson (BrianL): Blahp segfaults appeared to be due to overloaded NFS server
  • CTSC (BrianL): Attended meeting to discuss HTCondor issues encountered
  • LBNL (BrianL): Strange Gridmanager errors showing up intermittently

OSG Release Team

3.3.29 Δ Both Δ 3.4.4 Δ Total Δ Status
3 +3 12 +12 3 +3 18 +18 Open
0 +0 8 +8 2 +2 10 +10 In Progress
0 +0 1 +1 0 +0 1 +1 Ready for Testing
0 +0 0 +0 0 +0 0 +0 Ready for Release
3 +3 21 +21 5 +5 29 +29 Total
  • Both
    • Update globus-gridftp-server-control to 5.2
  • 3.4.4
  • 3.3.29

Discussions

OSG Investigations Team

Last Week

  • First step of creating ES snapshots for backup up GRACC - Ongoing.
  • Fixed explosion of "Fake" sites in records. Also fixed incorrect user VOs. Changing indexes to relefect the changes.

This Week

  • Initiate backups of ES snapshots
  • GRACC-ITB work
  • Start indexing GOC server status in GRACC ES
  • Stash Writeback hack-a-thon on Tuesday & Wednesday (Derek, Brian, Lincoln, Marian)

Ongoing

Discussions

Starting to setup time for Kibana walkthrough for Glidein Logs.